OpenStudy (anonymous):

Help Please. Find the Derivative of y=2xcosx-2sinx and apply chain rule. Can you show the process of how to solve it?

OpenStudy (anonymous):

if it is \[f(x)=2x\cos(x)-2\sin(x)\] then the chain rule does not apply only need the product rule

OpenStudy (anonymous):

so there is no chain rule at all?

OpenStudy (anonymous):

no because you have a product \(2x\sin(x)\) not a composite function

OpenStudy (anonymous):

Can you show me how to solve?